OpenStudy (naudiapie):

What is the equation of a circle that passes through point J(6, 4) and has center K (1, -8).

OpenStudy (naudiapie):

please help!!! :o

OpenStudy (mertsj):

Use the distance formula to calculate the length of the radius.

OpenStudy (naudiapie):

um can you show me? like give me an example of how i would set it up and everything

OpenStudy (naudiapie):

i don't know :(

OpenStudy (mertsj):

What is the distance formula?

OpenStudy (naudiapie):

D= sr(x2-x1)^2 + (y2-y1)^2

OpenStudy (mertsj):

Now plug in: the two points are: (6,4) and (1,-8)

OpenStudy (mertsj):

So x2=8 x1=6 y2=-8 y1=4

OpenStudy (naudiapie):

x= 8? wouldn't it be 1 though?

OpenStudy (naudiapie):

where did you get 2 8's at, ugh, this is confusing well i did 6-1 = 5 and 4 - 8 = -4 . ... so 5 - 4 = 1 , is 1 right?

OpenStudy (mertsj):

Sorry. x2=1

OpenStudy (naudiapie):

oh wait i forget to square them

OpenStudy (naudiapie):

so it would be like 20 - 16?

OpenStudy (naudiapie):

which is 4 and the square root of 4 is 2 ... idk :(

OpenStudy (mertsj):



OpenStudy (mertsj):

\[d=\sqrt{(1-6)^2+(-8-4)^2}=\sqrt{(-5)^2+(-12)^2}=\]

OpenStudy (mertsj):

Now you finish it.

OpenStudy (naudiapie):

169

OpenStudy (naudiapie):

13

OpenStudy (naudiapie):

thank you

OpenStudy (mertsj):

Can you finish the problem now that you know the radius is 13?

OpenStudy (naudiapie):

yes thank you lots :o

OpenStudy (mertsj):

yw